OpenStudy (anonymous):

Solve for x. 3x^2=-11-6

OpenStudy (anonymous):

\[3x^2=-17\] \[x^2=\frac{-17}{3}\] \[x=\pm \sqrt{\frac{-17}{3}}\] \[x=\pm (i \sqrt{\frac{17}{3}})\]